OpenStudy (anonymous):

If 1/2^m+1/2^m=1/2^x, then x expressed in terms of m is?

OpenStudy (kira_yamato):

Do you mean \[\frac{1}{2^m +1} + \frac{1}{2^m} = \frac{1}{2^x}\]?

OpenStudy (anonymous):

x=m+1

OpenStudy (anonymous):

the answer provided in the book is m-1 but I can't seem to figure out how to get there?

OpenStudy (anonymous):

@ Kira...eliminate the +1 in the first set..and thats what I meant

OpenStudy (anonymous):

I think it's still x=m+1